A particle of mass m moves under action of force F = – kr (k is positive constant). Find r(t) and v(t). (Boundary conditions r(0) = ro, v(0) = vo, at t=0). 


1
Expert's answer
2021-06-17T10:10:11-0400

Given,

Mass of the particle = m

Action of force "(F)=-kr"

"a=\frac{F}{m}"

Now,

"a=\frac{-kr}{m}"

We know that acceleration "a=\frac{dv}{dt}=\frac{dv dr}{dr dt}=v\frac{dv}{dr}"

Now, substituting the values,

"\Rightarrow \frac{vdv}{dr}=\frac{-kr}{m}"


"\Rightarrow vdv= \frac{-kr}{m}dr"

Now, taking the integration of both side of the given equation,

"\Rightarrow \int_ {v_1=0}^{v_2=v} vdv = \int_{r_1=0}^{r_2=r} \frac{-kr}{m}dr"


"\Rightarrow \frac{v^2}{2}=\frac{-kr^2}{2m}"